# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v6

@@ 5 Theoretical Foundation
We will introduce some new terms to help comprehend this section. The partitions we discussed earlier will now be formally referred to as containers. Additionally, we shall discuss the decreasing sequence property and its significance. After that, we will discuss interior, exterior, and back arcs. We will show the importance of these arcs with some foundational lemmas. The GLOVER data structure will then be used to wrap up.
@@
The formula $C_{i}=\{u\in V(G)\text{.s.t. }A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and }dist(v_{0},u)=i,dist(v_{0},v)=i\}$, where $V(G)$ denotes the graph $G$’s vertex set and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
Similar to this, Daamouch employed induced subgraphs in his research on $m$-free graphs [7]. The neighboring containers (see image), which will be crucial in the proofs, are the main distinction here.
We will investigate the structure of oriented graphs from two angles by using the idea of containers. Alternatively, we will investigate it from the viewpoint of containers or the minimum degree node.
  • Containers are seen based on their distance from the minimum degree node.
  • Nodes are seen based on their position relative to containers. They are either before a container, in a container, or after a container.
Specifically, the case where $d^{+}(v_{0})=3$ and Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve. A graph $G$ was created in which the out-degree of $v_{0}$ does not double in $G^{2}$. A simple graph of degree three corresponds to the lowest cycle that can exist in an undirected graph. Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the nodes adjacent to $v_{0}$. A closer examination of this case reveals that the out-degree of a node that is an out-neighbor of $v_{0}$ is (at least) twice in this situation. This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
However, we will need to define new concepts in order to proceed with this line of thinking. To prevent a node’s out-degree from doubling, we must first find a term. The decreasing sequence property is attributed to a node in $G$ whose out-degree does not double in $G^{2}$. There are instances where $v_{0}$ lacks its out-degree double in $G^{2}$, as we have seen. We’ll investigate which other nodes can possess the decreasing sequence property and which network features are necessary for this to occur.
@@
Two new ideas have been presented thus far: the decreasing sequence property (DSP) and the minimal degree node. A crucial starting point for our approach and degree values is a minimal degree node, which is the node with the fewest connections. It acts as the anchor by being chosen as the beginning point. The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node. It is serving as a stimulant.
In contrast, the DSP introduces constraints on connectivity. This concept is akin to a bottleneck. Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
@@
Let $G$ be a graph with a minimum-degree-node $v_{0}$. Construct a series of containers $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{k}$ ordered by distance relative to $v_{0}$. Each container $C_{i}$ is a set of nodes at distance $i$ from $v_{0}$. This imposes a total order on the nodes in $G$. This means that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
@@
While imposing order on our graph is a solid first step, the containers do more than just enforce overall order. In essence, they enable group comparisons between nodes that are equally spaced from a reference node, $v_{0}$. A typical total order, such as lexicographical order, does not take this into consideration. This ”containerized ordering” allows us to differentiate nodes based on particular adjacency qualities inside each distance group in addition to their distance from $v_{0}$.
@@ A Level Total Order (LTO) on a set $S$ can be defined as follows:
  • Leveled Structure: The elements of $S$ are partitioned into a leveling of containers. $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{n}$ Each container itself forms a subset of $S$. This maintains a total order. The containers are sequentially ordered relative to each other. This establishes a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
  • Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$. This avoids trivial total orderings where all elements are isolated in singleton containers.
@@
  • Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves.
@@
The notion of a non-triviality requirement is to distinguish this ordering from other traditional total orderings on graphs.
@@
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$. In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed. We will now give terms to these concepts.
@@
In particular, we saw things about interior neighbors in 4 (Minimum Out-Degree 3 with Neighbors 1). The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1, and we see that when all the neighbors in that subgraph had an out-degree of 1, $v_{0}$’s degree did not double.
@@
We contrast interior neighbors with exterior neighbors, which reveal how a container interacts with adjacent containers. This interaction is again critical for the Seymour Second Neighborhood Conjecture, as the nodes within the current container depend on the connections to the next container.
@@
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. In order to do so, though, we need to consider the option of back-arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to the minimum degree node. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes.
@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$.
@@
To prove that any oriented graph can be represented in a GLOVER, we’ll need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties.
@@
Assume we have an oriented graph $G$. Assuming that there are no back arcs in $G$, this oriented graph can be represented by a GLOVER.
@@
To establish a Graph Level Order, we first need a minimum degree node. We can obtain this by taking the minimum value on the degree vector of nodes. We will call that in the node $v_{0}$.
Now that we know the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$.
Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes whose distance from the node $v_{0}$ is equal. These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
  • Total Order on Elements within $G$ Because nodes are partitioned based on their distance from $v_{0}$, we have a total order on the elements of $G$. For any two nodes $u$ and $v$, this total order is determined by their containers.
  • Leveled Structure with Containers - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ form ranked levels, each corresponding to nodes at a specific distance from $v_{0}$.
  • Non-triviality Requirement - Given that $v_{0}$ is a minimum degree node, there is at least one container (specifically, $C_{1}$) that has more than one node.
  • Comparability within Containers - For any two nodes $u$ and $v$ in the same container $C_{i}$, the order can be determined by their relationship to $v_{0}$.
@@
    • Nodes $u$ and $v$ of the same distance from $v_{0}$ would be in the same container and thus be interior neighbors.
@@
What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers. This ordering is also ordering the nodes within the containers, so the GLOVER data structure is a multi-level total order.
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each container maintains its place in the total ordering.
Next, we will show one of the other benefits of this data structure in mathematical induction via well ordering.
@@
GLOVER Supports Induction
In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
@@
We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering.
A set is considered well-ordered if every non-empty subset of that set has a least element.
A discrete total order is a total order where each element has a distinct ”next” element in the order.
Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a ”smallest” element according to the order. This fulfills the condition of being well-ordered.
Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set.
@@
Interior Degree at least 1
Suppose we have an oriented graph $G$, represented by a GLOVER. Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property. In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $N^{+}(v_{0})$ must have an interior degree of at least 1. That is, for $x\in N^{+}(v_{0})$,
@@
By definition, a node satisfies the decreasing sequence property if
$|N^{+}(x)|>N^{++}(x)|$
For $v_{0}$, the second neighbors are the nodes reachable by following an arc from $v_{0}$ to a neighbor $x$, and then from $x$ to another node $z$, i.e., $z\in N^{++}(v_{0})$.
Since $v_{0}$ is the minimum out-degree node, we know that $d^{+}(v_{0})$ is smaller than or equal to the out-degree of any other node. Let $x\in N^{+}(v_{0})$.
If $x$ is connected to $d^{+}(v_{0})$ or more nodes via exterior arcs, this would cause $v_{0}$’s degree to double through $x$’s neighbors, which are $v_{0}$’s second neighbors. If $x$ connects to $v_{0}$ through a back arc, again this would cause $v_{0}$’s degree to double through $x$’s first neighbors.
To avoid this doubling, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
@@
This argument holds for any neighbor $x\in N^{+}(v_{0})$. This ensures that every node in $N^{+}(v_{0})$ has an interior degree of at least 1.
Thus, for all $x\in N^{+}(v_{0})$, $|int(v_{0},x)|\geq 1$.
Lemma 5.4 (Interior Degree at least 1) is important because we can see some structural properties in general oriented graphs that may coincide with degree doubling. In particular, we see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree to double.
Example 5.4.
0
@@
Exterior Less than Degree
Suppose we have an oriented graph $G$, represented by a GLOVER. In an attempt to prevent a degree-doubling node, if $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i}),|\text{ext}(x,y)|<d^{+}(x)$.
@@
Assume, for the sake of contradiction, that there exists a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$.
By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$
Thus, $|\text{ext}(x,y)|\subseteq N^{++}(x)$, then:
@@
suggesting that the degree of outwardness of $x$ would at least double on graph $G^{2}$. This is a contradiction to the assumption that $x$ satisfies the decreasing sequence property.
Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|\text{ext}(x,y)|$ must be strictly less than $d^{+}(x)$.
@@
Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an exhaustive search solution with the algorithm we are developing.
@@
Weak Decreasing Exteriors Lemma
Suppose that $G$ is an oriented graph represented by a GLOVER, and that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
@@
    The interior degree of $y$ with respect to $x$, $|int(x,y)|$, doubles in the subgraph $G[N^{+}(x)]^{2}$.
@@
We begin by recalling the conditions for a node to have the decreasing sequence property (DSP). A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$.
Now, assume the conditions of the lemma hold:
@@
    The interior degree $|int(x,y)|$, doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increase significantly in $G[N^{+}(x)]$.
@@
    For all $z\in\text{ext}(x,y)$, $\text{ext}(y,z)\cap N^{+}(x)=\emptyset$, ensuring that no back arcs exist.
Now, suppose for contradiction that there exists $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$.
The node $y$ has a degree that can be broken down into two components, its interior degrees and its exterior degrees. By the assumption that $y$’s interior degree doubles, if $y$’s exterior degree were to double, it would lead to a contradiction.
If there exists a node $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$, then $y$’s out-degree will be impacted by $\text{ext}(x,y)$ and $\text{ext}(y,z)$.
@@
Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$.
The first thing to notice in this proof is the assumptions. We assume that there are no back arcs and that a node’s interior degree doubles. Both of these assumptions are important.
First, we will talk about the lack of back arcs. We have already talked about how back arcs can be a hindrance to our algorithm overall. This is one of the main points where they can be a hindrance. In particular, there exists a way to double count nodes as both neighbors and second neighbors if back arcs exist. We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
Secondly, we assumed that the interior degree of a node doubled. This is similar to the notion of solving the containers separately, then joining them together again to solve the main problem. We are assuming that this node $y$ was in another container and its interior degree doubled in that container. In doing so, $y$ is now in a better position to have its overall degree double, only needing to have its exterior degree double.
The more popular representations of graphs are adjacency matrices, arraylists and edge lists. These suffer from the same limitations of graphs as mentioned above, They scatter the nodes along the two-dimensional plane without attempts to group into an ordering. The Graph Level Order (GLOVER) improves this by partitioning the nodes in a reasonable way. For the SSNC, that reasonable way was the distance metric. In order to represent graphs in a GLOVER we would use any markup language that can encode graphs. There needs to be three types of data- containers, nodes, and arcs. The containers will be given ids, the nodes will then be assigned to their unique containers, and arcs will be determined to be interior, exterior, or back, depending on the containers of the endpoints of the arcs.
This determination of the number of containers and the assignment of nodes to containers does not need to be declared beforehand. Instead, just as we have been talking about the distance of a node from a minimum degree node $v_{0}$ as the determination of container assignment, we would have the same or a similar function to assign nodes to containers.

### v7

@@ 5 Theoretical Foundation
We will introduce some new terms to help with the rest of this paper. The partitions we discussed earlier will now be formally referred to as containers. Additionally, we shall discuss the decreasing sequence property and its significance. After that, we will discuss interior, exterior, and back arcs. We will show the importance of these arcs with some foundational lemmas. The GLOVER data structure will then be used to wrap up.
@@
The formula
$C_{i}=\{u\in V(G)\text{.s.t. }A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and }dist(v_{0},u)=i,dist(v_{0},v)=i\}$, where
$V(G)$ denotes the graph $G$’s vertex set and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
Similar to this, Daamouch employed induced subgraphs in his research on $m$-free graphs [7]. The neighboring containers (see example 5.1), which will be crucial in the proofs, are the main distinction here.
We will investigate the structure of oriented graphs from two angles by using the idea of containers. Alternatively, we will investigate it from the viewpoint of containers or the minimum degree node. Containers are seen based on their distance from the minimum degree node. Nodes are seen based on their position relative to containers. They are either before a container, in a container, or after a container.
Specifically, the case where $d^{+}(v_{0})=3$ and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve. A graph $G$ was constructed in which the out-degree of $v_{0}$ does not double in $G^{2}$. A simple graph of degree three corresponds to the smallest cycle that can exist in an undirected graph. Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the nodes adjacent to $v_{0}$. If we look at this situation further, we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double. This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
However, we will need to define new concepts to proceed with this line of reasoning. To prevent a node’s out-degree from doubling, we must first define a term. The decreasing sequence property is attributed to a node in $G$ whose out-degree does not double in $G^{2}$. There are instances where $v_{0}$ lacks its out-degree double in $G^{2}$, as we have seen. We’ll investigate which other nodes can possess the decreasing sequence property and which network features are necessary for this to occur.
@@
Two competing new ideas have been presented, the decreasing sequence property (DSP) and the minimal degree node. A crucial starting point for our approach and degree values is a minimal degree node. It acts as the anchor by being chosen as the beginning point. The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node. It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
In contrast, the DSP introduces constraints on connectivity. This concept is akin to a bottleneck. A bottleneck in graph theory represents a point that limits flow or connectivity within the network, a role precisely filled by the DSP. Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
@@
Let $G$ be a graph with a minimum-degree node $v_{0}$. Construct a series of containers $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{k}$ ordered by distance relative to $v_{0}$. Each container $C_{i}$ is a set of nodes at distance $i$ from $v_{0}$. This imposes a total order on the nodes in $G$. This means that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
@@
While imposing order on our graph is a solid first step, the containers do more than just enforce overall order. In essence, they allow for the grouping of comparisons between nodes that are equally spaced from a reference node, $v_{0}$. A typical total order, like lexicographical order, does not take this into consideration. This ”containerized ordering” allows us to differentiate nodes based on particular adjacency qualities inside each distance group in addition to their distance from $v_{0}$.
@@ A Level Total Order (LTO) on a set $S$ can be defined as follows:
  • Leveled Structure:
    • The elements of $S$ are partitioned into a leveling of containers. $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{n}$
    • Each container itself forms a subset of $S$.
    • The containers are sequentially ordered relative to each other.
@@
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$. In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed. We will now give terms to these concepts.
@@
In particular, we saw things about interior neighbors in 4.3 (Minimum Out-Degree 3 with Neighbors 1). The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1, and we see that when all the neighbors in that subgraph had an out-degree of 1, $v_{0}$’s degree did not double.
@@
We contrast interior neighbors with exterior neighbors, which reveals how a container interacts with adjacent containers. This interaction is again critical for the SSNC, as the nodes within the current container depend on the connections to the next container.
@@
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. To do so, though, we need to consider the option of back arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to the minimum degree node. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes.
@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
Next we will prove that any oriented graph can be represented in a GLOVER. We will need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties. This includes maintaining a total order of the containers, allowing for the comparison of any two nodes, establishing a total order of those nodes within the containers, and establishing interior and exterior neighbors.
@@
Given an oriented graph $G$. Assuming that there are no back arcs in $G$, this oriented graph can be represented by a GLOVER data structure.
@@
To establish a graph level order, we first need a minimum degree node. So we take the minimum value on the degree vector of nodes. We will call that node $v_{0}$. Knowing the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$. Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes at distance $i$ from $v_{0}$. These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
  • Total Order on Elements within $G$ Nodes are partitioned based on their distance from $v_{0}$. This is a total order on the nodes of $G$. For any two nodes $u$ and $v$, this total order is determined by their containers.
  • Leveled Structure with Containers - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ each correspond to nodes at a specific distance from $v_{0}$.
  • Comparability within Containers - For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric to distinguish $u$ from $v$.
@@
    • Nodes $u$ and $v$ of the same distance from $v_{0}$ are in the same container and thus are interior neighbors.
@@
What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers. This ordering is more than just representative, though. The containers help to organize the nodes. What this suggests is that each partition, or set of nodes, is being totally ordered as we move away from the chosen minimum degree node. This is a clear distinction from standard graph theory. Graph nodes are generally thought of as able to be placed anywhere in 2-dimensional space. However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from the chosen minimum degree node.
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each container maintains its place in the total ordering. However, it is a bold assumption to claim that oriented graphs have no back arcs. Oriented graphs come in various types; surely some exist with back arcs.
This is an important claim to investigate, and the next section will look into it in depth. We will prove that these back arcs do not hurt the algorithm. Instead, they only lead to the faster path to a degree-doubling node. In addition, there will no longer be a need to represent the oriented graph in a GLOVER, as we will have already found the degree-doubling node.
Next, we will show one of the other benefits of this data structure in mathematical induction via well ordering. Mathematical induction is a proof technique that uses inference to prove statements. Remember that a well ordering is a concept in discrete mathematics where a total ordering on a set $S$ has the property that every non-empty subset of $S$ has a least element.
@@
GLOVER Supports Induction In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
@@
We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering. A set is considered well-ordered if every non-empty subset of that set has a least element. A discrete total order is a total order where each element has a distinct ”next” element in the order. Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a ”smallest” element according to the order. This fulfills the condition of being well-ordered. Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set. ∎
This becomes an important concept because now, in addition to the nodes, we also have the containers as subsets of $S$. These containers are totally ordered and thus can also be used as a means of mathematical induction. We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from the chosen minimum degree node $v_{0}$.
@@
Interior Degree at least 1 Suppose we have an oriented graph $G$, represented by a GLOVER. Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property. In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $C_{1}$ must have an interior degree of at least 1. That is, for $x\in C_{1}$,
@@
By definition, a node satisfies the decreasing sequence property if $|N^{+}(x)|>N^{++}(x)|$ If $x$ has $d^{+}(v_{0})$ exterior arcs, this would make $v_{0}$’s degree double because $x$’s neighbors are $v_{0}$’s second neighbors. To avoid this doubling, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
@@
This is true for any neighbor $x\in C_{1}$. This ensures that every node in $C_{1}$ has an interior degree of at least 1. Thus, for all $x\in C_{1}$, $|int(v_{0},x)|\geq 1$. ∎
Lemma 5.4 (Interior Degree at least 1) is important because we can see some structural properties in generally oriented graphs that may coincide with degree doubling. We see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree doubled.
@@
Exterior Less than Degree Suppose we have an oriented graph $G$, represented by a GLOVER. In an attempt to prevent a degree-doubling node, if $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
@@
Assume, for the sake of contradiction, that there is a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$. By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$. Thus, $|\text{ext}(x,y)|\subseteq N^{++}(x)$. Then:
@@
This means that the out-degree of $x$ at least doubles on the graph $G^{2}$. This contradicts the assumption that $x$ has the decreasing sequence property.
This must be true for all $x$ in $G$, so we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|\text{ext}(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
@@
Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an global search solution with the algorithm we are developing.
@@
Weak Decreasing Exteriors Lemma Suppose that $G$ is an oriented graph represented by a GLOVER. Suppose further that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
@@
    The interior degree of $y$ with respect to $x$, $|int(x,y)|$, doubles in the induced subgraph $G[N^{+}(x)]^{2}$.
@@
We begin by recalling the conditions for a node to have the decreasing sequence property (DSP) because it is very important for this proof. A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$. Now, assume the conditions of the lemma hold:
@@
    The interior degree $int(x,y)$ doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increase significantly in $G[N^{+}(x)]$.
@@
    For all $z\in\text{ext}(x,y)$, $\text{ext}(y,z)\cap N^{+}(x)=\emptyset$; no back arcs exist.
Suppose for contradiction that there exists $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$. The node $y$ has a degree that can be broken down into two components, its interior degrees and its exterior degrees. By the assumption that $y$’s interior degree doubles, if $y$’s exterior degree were to double, it would lead to a contradiction. If there exists a node $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$, then $y$’s out-degree will be impacted by $\text{ext}(x,y)$ and $\text{ext}(y,z)$.
@@
Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$. ∎
The first thing to notice in this proof is the assumptions. We assume that there are no back arcs from $x$’s second neighbors to $x$’s first neighbors and that a $y$’s interior degree doubles. Both of these assumptions are important.
First, we will talk about the lack of back arcs. We have already talked about how back arcs can be a hindrance to our algorithm overall. This is one of the main points where they can be a hindrance. There exists a way to double count nodes as both first neighbors and second neighbors if back arcs exist. We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
Secondly, we assumed that the interior degree of a node doubled. This is like the notion of solving the containers separately, then joining them together again to solve the main problem. We are assuming that this node $y$ was in another container and its interior degree doubled in that container. In doing so, $y$ is now in a better position to have its overall degree double, only needing to have its exterior degree double.
Notice the wording of interior degree doubling. That is a separate version of the SSNC, but localized to the containers. The conjecture in this case is can we find a node in each container whose degree doubles. If so, then it gives light to this divide-and-conquer approach, where we partition the containers, solve them individually, and then combine them again for this lemma 5.6 (Weak Decreasing Exteriors Lemma).
The more popular representations of graphs are adjacency matrices, arraylists, and edge lists. These suffer from the same limitations of graphs as mentioned above; they scatter the nodes along the two-dimensional plane without attempts to group them into an ordering. The Graph Level Order (GLOVER) improves this by partitioning the nodes in a reasonable way. For the SSNC, that reasonable way was the distance metric. To represent graphs in a GLOVER, we can use any data serialization that can encode graphs, including JSON, YAML, and XML. There needs to be three types of data: containers, nodes, and arcs. The containers will be given IDs. The nodes will then be assigned to their unique containers. Arcs will also be determined to be interior, exterior, or back, depending on the endpoints of the nodes of the arcs.
The determination of the number of containers and the assignment of nodes to containers also do not need to be declared beforehand. Instead, just as the data structure uses the minimum function to find a node representing the minimum degree node $v_{0}$. This same function will return a value $\delta=d^{+}(v_{0})$, and $\delta$ will represent the maximum number of containers. We would have a similar function to assign nodes to containers.